# Long-term stable nonlinear evolutions of ultracompact black-hole mimickers

### v1

@@ VII Concluding remarks
We have performed a series of fully nonlinear evolutions of the Einstein-Klein-Gordon equations governing the dynamics of solitonic boson stars in the thin-shell regime in 3+1 dimensions imposing spherical, axi-, octant or no symmetry. A subset of the models under consideration are ultracompact, exhibiting a pair of light rings, one of which is stable. We have evolved these objects for $\gtrsim$1000 light crossing times, finding no evidence of an instability due to the appearance of the stable light ring, as previously conjectured Keir (2016); Cardoso et al. (2014); Cunha et al. (2023). We observe excellent agreement between the axisymmetric/$3D$ evolutions and those in spherical symmetry, suggesting that the inclusion of nonspherical degrees of freedom does not significantly affect the stability of these models. Taking also into account the agreement of our numerical results with perturbative predictions for stability and fundamental radial oscillation frequencies, we interpret our findings as strong numerical evidence in favor of the stability of this class of BH mimickers.
@@
As a cautionary note, we add that poor choices of numerical ingredients, in particular the constraint damping parameter $\kappa_{1}$ of the CCZ4 formulation, can result in seemingly plausible instabilities of our thin-shell BSs. With $M\kappa_{1}\gg 0.1$, for example, both models S06A044 and S08A06 collapse to a BH; the erroneous nature of these simulations reveals itself only through large and growing constraint violations and complete lack of convergence.
@@
If the presence of the stable light ring is not sufficient to destroy these solitonic boson stars, this family of spherically symmetric models, as well as similar UCOs, can be considered as viable black hole mimickers again. The thin-shell solutions under consideration here can mimick the phenomenology of black holes in several different ways, as for example in the GW emission from binaries Evstafyeva et al. (2024) or the creation of BH like shadows in VLBI images, Rosa et al. (2023). Figure 6 illustrates how the S06A044 model could resemble a Schwarzschild BH with the same ADM mass in the presence of an equatorial, optically thin accretion disk. The image of the BS clearly displays additional structure towards the center of the image, which is a result of the absence of a horizon as well as the pair of light rings. The outermost direct emission band, as well as the first lensed photon ring are nearly indistinguishable for both images.
@@
Figure 6: Side-by-side comparison of ray-traced images of a (left) Schwarzschild BH and (right) the S06_A044 model of the same ADM mass. The color scale indicates the observed brightness intensity, in arbitrary units, from a simple geometrically thin disk located in the equatorial plane. Both panels also provide a zoom-in of the photon ring region, exhibiting their fractal structure. Images are obtained with FOORT Mayerson (2024); Mayerson et al. - see App. E.
Acknowledgements.
@@
G.A.M. is supported by the Cambridge Trust at the University of Cambridge. S.S. is supported by the Centre for Doctoral Training at the University of Cambridge funded through STFC. TE acknowledges Perimeter Institute for Theoretical Physics, supported by the Government of Canada through the Department of Innovation, Science and Economic Development and by the Province of Ontario through the Ministry of Colleges and Universities. This work has been supported by STFC Research Grant No. ST/V005669/1. We acknowledge support by the NSF Grant Nos. PHY-090003, PHY-1626190 and PHY-2110594, DiRAC projects ACTP284 and ACTP238, STFC capital Grants Nos. ST/P002307/1, ST/R002452/1, ST/I006285/1 and ST/V005618/1, STFC operations Grant No. ST/R00689X/1. Computations were done on the CSD3 and Fawcett (Cambridge), Cosma (Durham), Hawk (Cardiff), CIT LIGO Lab (Pasadena), Stampede2 (TACC) and Expanse (SDSC) clusters.
Appendix A Convergence Analysis
@@
Light rings: We check convergence of the computed light ring (LR) radii for the S08A06 model by comparing results obtained at resolutions $\mu h\in\{1/12,1/16,1/20\}$ and display in Fig. 7 the resulting areal radii of the stable LR as functions of time. At low resolution the LR radius exhibits a linear drift, much like what was observed for the central amplitude and Kretschmann scalar; cf. Fig. 2. Again, this drift converges away at 5th order, and the Richardson extrapolated values, obtained from the two highest-resolution runs and displayed in Fig. 8, oscillate around constant values.
@@
Figure 7: Evolution of the areal radius of the stable LR for the S08A06 model obtained with the E3D code using different resolutions indicated in the legend. The differences between the curves yield convergence at 5th order.
@@
Figure 8: Richardson extrapolation (5th order) of the LR areal radii of the S08A06 model based on the runs at resolutions $\mu h=1/16,1/20$. The bottom panel shows the evolution of the potential difference between the extrema of the AEP.
@@
2D code: We next discuss the convergence results for the S08A06 model obtained in axisymmetric evolutions with the Axi code. Using $\mu h=1/8$, $\mu h=1/12$ and $\mu h=1/16$ on the innermost refinement level for the low, medium and high resolutions, respectively, we obtain convergence between fourth and fifth order for the scalar field amplitude; cf. top panel of Fig. 9. In the bottom panels of the figure, we also monitor the volume-weighted $L_{2}$ norms of the constraint violations. We find first-order convergence in the Hamiltonian constraint at very early times which gradually transitions to fourth order at $\mu t\approx 20000$ and persists throughout the remainder of the simulation. The momentum constraint on the other hand rapidly reaches convergence between fourth and fifth order.
@@
Figure 9: Top (bottom) panels illustrate convergence results obtained from Axi simulations of the S08A06 model for the scalar field amplitude (volume-weighted $L_{2}$ norms of the Hamiltonian and momentum constraints). In the bottom panels we display the constraint violations for the low and medium resolutions and amplify the latter by factors $Q_{n}$ as expected for $n^{\rm th}$-order convergence.
Appendix B Additional results for $\sigma_{0}=0.06$
@@
In this section we present results for the $\sigma_{0}=0.06$ family of models, supplementing those presented for $\sigma_{0}=0.08$ above. In Fig. 10, we compare the power spectrum of central scalar-amplitude oscillations to the radial oscillation frequency, computed with perturbation theory. In this case the larger error in our initial data causes the system to migrate early on to a BS model with slightly smaller central scalar amplitude than the initial value $A_{0}=0.044$. Following Ref. Kain (2021) we take this effect into account by computing the radial frequency for a BS model with an amplitude averaged over several oscillation cycles, $\langle A_{0}\rangle=0.04394$. Once again, we find very good agreement between results from the perturbative analysis and the codes enforcing different symmetry assumptions.
Figure 10: Power spectrum of $\Delta A_{0}(t)$ for S06A044; the red line marks the fundamental radial mode frequency $\chi_{0}$. Units on the vertical axis are arbitrary.
@@
In Fig. 11, we show timescales of instability for a sample of BS models with $\sigma_{0}=0.06$. As before, we define the instability timescale as the time at which the absolute change in the central amplitude $|\Delta A_{0}(t)|$ exceeds $10\%$ of the initial central amplitude $\Delta A_{0}(0)$. We find that the predictions of the radial perturbation theory are borne out by all codes regardless of symmetry assumptions. Models on the first unstable branch either migrate or disperse, and models on the second unstable branch always collapse to a BH. We emphasize that the actual timescales depend strongly on factors such as resolution, gauge, evolution method and parameter choices, mesh refinement, and other code differences; they should therefore only be interpreted as having a relative meaning. Note also that for some values of the central amplitude, multiple BS models exist; where ambiguous it is always the two lower-mass models which are stable, in accordance with the predictions of linear theory.
Figure 11: Instability timescales as a function of $A_{0}$ for solitonic BSs with $\sigma_{0}=0.06.$ The vertical lines mark the transition between stable and unstable branches as predicted by linear theory. Marker style indicates dynamical fate while colour indicates code used: black is Sph, blue Axi, orange E3D. BS models whose time evolution show no sign of instability are represented with arrows at the top of the figure.
Appendix C Radial Oscillation Frequencies
@@
In this section we discuss in detail the method used to compute fundamental radial oscillation frequencies for solitonic boson stars (BSs), presenting results across both stable branches for a variety of $\sigma_{0}$-values. In this section, we use overdots to denote time derivatives and primes to denote radial derivatives.
We begin with a spherically symmetric metric ansatz
$\mathrm{d}s^{2}=-\alpha^{2}(t,r)\mathrm{d}t^{2}+X^{2}(t,r)\mathrm{d}r^{2}+r^{2}\mathrm{d}\Omega_{2}^{2}.$ (4)
The Einstein field equations then reduce to the following, where we write $\varphi(t,r)=\varphi_{1}+\mathrm{i}\mkern 1.0mu\varphi_{2}$ for real $\varphi_{1}$, $\varphi_{2}$,
$\frac{X^{\prime}}{X}=\,$ $\frac{1-X^{2}}{2r}+2\pi rX^{2}\left(\frac{|\dot{\varphi}|^{2}}{\alpha^{2}}+\frac{|\varphi^{\prime}|^{2}}{X^{2}}+V\right),$ (5)
$\frac{\dot{X}}{X}=\,$ $4\pi r\left(\dot{\varphi}_{1}\varphi_{1}^{\prime}+\dot{\varphi}_{2}\varphi_{2}^{\prime}\right),$ (6)
$\frac{\alpha^{\prime}}{\alpha}=\,$ $-\frac{1-X^{2}}{2r}+2\pi rX^{2}\left(\frac{|\dot{\varphi}|^{2}}{\alpha^{2}}+\frac{|\varphi^{\prime}|^{2}}{X^{2}}-V\right),$ (7)
$\alpha^{\prime\prime}=\,$ $\frac{X\ddot{X}}{\alpha}-\frac{X\dot{X}\dot{\alpha}}{\alpha^{2}}+\frac{X^{\prime}\alpha^{\prime}}{X}-\frac{\alpha^{\prime}}{r}+\frac{\alpha X^{\prime}}{rX}$ (8)
@@
Because of the Bianchi identities these equations are not independent. At first glance, one may be tempted to simply discard the final equation as a direct consequence of the former three, but later on Eq. (8) turns out to be useful for converting the pulsation equations into a system of two second-order partial differential equations (PDEs) in two variables. The Einstein equations are complemented by the Klein-Gordon equation
$\varphi_{i}^{\prime\prime}=\left(\frac{X^{\prime}}{X}-\frac{\alpha^{\prime}}{\alpha}-\frac{2}{r}\right)\varphi_{i}^{\prime}+\frac{X^{2}\dot{\alpha}+X\alpha\dot{X}}{\alpha^{3}}\dot{\varphi}_{i}+\frac{X^{2}}{\alpha^{2}}\ddot{\varphi}_{i}$ (9)
where $i\in\{1,2\}$. Together, Eqs. (5)-(9) completely characterize our Einstein-Klein-Gordon system.
We now wish to expand around a background BS spacetime, particularizing at this point to the solitonic potential. For this purpose we introduce $\delta\psi_{1}$, $\delta\psi_{2}$, $\delta\nu$ and $\delta\lambda$ as perturbation functions of $(t,r)$ according to
$\varphi$ $=e^{\mathrm{i}\mkern 1.0mu\omega t}\left(\psi_{1}+\mathrm{i}\mkern 1.0mu\psi_{2}\right)\,,$
$\psi_{1}$ $=A_{0}(r)(1+\delta\psi_{1}),\;\psi_{2}=A_{0}(r)\delta\psi_{2},\;$
$\alpha$ $=\alpha_{0}(r)\left(1+\frac{1}{2}\delta\nu\right),\ X=X_{0}(r)\left(1+\frac{1}{2}\delta\lambda\right).$ (10)
At zeroth order in the perturbations, this yields the familiar ordinary-differential-equation (ODE) system for a stationary BS,
$A_{0}^{\prime\prime}$ $=-\frac{\omega^{2}X_{0}^{2}}{\alpha_{0}^{2}}A_{0}+A_{0}^{\prime}\left(\frac{X^{\prime}_{0}}{X_{0}}-\frac{\alpha^{\prime}_{0}}{\alpha_{0}}-\frac{2}{r}\right)+X_{0}^{2}A_{0}V^{\prime}(A_{0}^{2}),$ (11)
$\frac{X_{0}^{\prime}}{X_{0}}$ $=\frac{1-X_{0}^{2}}{2r}+2\pi r\left[(A_{0}^{\prime})^{2}+\frac{X_{0}^{2}}{\alpha_{0}^{2}}\omega^{2}A_{0}^{2}+X_{0}^{2}V(A_{0}^{2})\right],$ (12)
$\frac{\alpha_{0}^{\prime}}{\alpha_{0}}$ $=\frac{X_{0}^{2}-1}{2r}+2\pi r\left[(A_{0}^{\prime})^{2}+\frac{X_{0}^{2}}{\alpha_{0}^{2}}\omega^{2}A_{0}^{2}-X_{0}^{2}V(A_{0}^{2})\right].$ (13)
At linear order, after simplifying background terms by using Eqs. (11), (12) we obtain
@@
$\delta\psi_{1}^{\prime\prime}=\,$ $X_{0}^{2}\left(1-\frac{\omega^{2}}{\alpha_{0}^{2}}-\frac{8A_{0}^{2}}{\sigma_{0}^{2}}+\frac{12A_{0}^{4}}{\sigma_{0}^{4}}\right)\delta\lambda+\frac{X_{0}^{2}}{\alpha_{0}^{2}}\delta\ddot{\psi}_{1}+\frac{16A_{0}^{2}X_{0}^{2}}{\sigma_{0}^{2}}\left(\frac{3A_{0}^{2}}{\sigma_{0}^{2}}-1\right)\delta\psi_{1}+\frac{A_{0}^{\prime}}{2A_{0}}\left(\delta\lambda^{\prime}+4\delta\psi_{1}^{\prime}-\delta\nu^{\prime}\right)$
$+\frac{\omega X_{0}^{2}}{\alpha_{0}^{2}}\left(2\delta\dot{\psi}_{2}-\omega\delta\nu\right)+\left(\frac{X_{0}^{\prime}}{X_{0}}-\frac{\alpha_{0}^{\prime}}{\alpha_{0}}-\frac{2}{r}\right)\delta\psi_{1}^{\prime}\,,$ (14)
@@
$\delta\lambda^{\prime}=\,$ $\left(\frac{2X_{0}^{\prime}}{X_{0}}-\frac{1}{r}-4\pi r(A_{0}^{\prime})^{2}\right)\delta\lambda+8\pi rX_{0}^{2}\left(A_{0}^{2}+\frac{(A_{0}^{\prime})^{2}}{X_{0}^{2}}+\frac{\omega^{2}A_{0}^{2}}{\alpha_{0}^{2}}+\frac{12A_{0}^{6}}{\sigma_{0}^{4}}-\frac{6A_{0}^{4}}{\sigma_{0}^{2}}\right)\delta\psi_{1}$
$+\frac{4\pi r\omega A_{0}^{2}X_{0}^{2}}{\alpha_{0}^{2}}\left(\omega\delta\nu+2\delta\dot{\psi}_{2}\right)+8\pi rA_{0}^{\prime}A_{0}\delta\psi_{1}^{\prime}.$ (15)
@@
$\delta\ddot{\lambda}=\,$ $\frac{2\alpha_{0}}{X_{0}^{2}}\left[\frac{\alpha_{0}X_{0}^{\prime}}{rX_{0}}+\alpha_{0}^{\prime}\left(\frac{X_{0}^{\prime}}{X_{0}}-\frac{1}{r}\right)-\alpha_{0}^{\prime\prime}-4\pi\alpha_{0}(A_{0}^{\prime})^{2}\right]\delta\lambda-\frac{16\pi\alpha_{0}^{2}}{\omega X_{0}}\left[\frac{(A_{0}^{\prime})^{2}}{X_{0}^{2}}+A_{0}^{2}-\frac{8A_{0}^{4}}{\sigma_{0}^{2}}+\frac{12A_{0}^{6}}{\sigma_{0}^{4}}-\omega^{2}A_{0}^{2}\right]\delta\psi_{1}$
@@
$+8\pi\omega A_{0}^{2}\left(\omega\delta\nu-2\delta\dot{\psi}_{2}\right)+\frac{1}{rX_{0}}\left(\frac{r\alpha_{0}^{2}\delta\nu^{\prime}}{X_{0}}\right)^{\prime}+\frac{16\pi\alpha_{0}^{2}A_{0}^{\prime}A_{0}}{X_{0}^{2}}\delta\psi_{1}^{\prime}-\frac{1}{rX_{0}^{2}}\left(\alpha_{0}^{2}+r\alpha_{0}\alpha_{0}^{\prime}\right)\delta\lambda^{\prime}.$ (16)
Finally, it is convenient to also linearize the difference between Eqs. (5) and (7), which leads to
@@
$\delta\lambda^{\prime}-\delta\nu^{\prime}=\,$ $2\left(\frac{X_{0}^{\prime}}{X_{0}}-\frac{\alpha_{0}^{\prime}}{\alpha_{0}}-\frac{1}{r}\right)$ (17)
$+16\pi rA_{0}^{2}X_{0}^{2}\left(1-\frac{8A_{0}^{2}}{\sigma_{0}^{2}}+\frac{12A_{0}^{4}}{\sigma_{0}^{4}}\right).$
@@
Next, we make a harmonic ansatz $\delta\psi_{1}(t,r)=e^{\mathrm{i}\mkern 1.0mu\chi t}f(r)$, $\delta\lambda(t,r)=e^{\mathrm{i}\mkern 1.0mu\chi t}g(r)$ with $\chi$ undetermined. The resulting system of two pulsation equations is self-adjoint, so that, in analogy to the BS frequencies $\omega$, the radial oscillation frequencies $\chi$ form an infinite discrete sequence $\chi_{0}^{2}<\chi_{1}^{2}<...$, where the solution with eigenvalue $\chi_{n}^{2}$ has $n$ zero crossings in $f$ and $g$. A negative $\chi_{0}^{2}$ corresponds to a radial instability, while $\chi_{0}^{2}>0$ guarantees (linear) radial stability.

### v2

@@ Abstract
Concluding remarks—We have performed a series of fully nonlinear evolutions of the Einstein-Klein-Gordon equations governing the dynamics of solitonic boson stars in the thin-shell regime in 3+1 dimensions imposing spherical, axi-, octant or no symmetry. A subset of the models under consideration are ultracompact, exhibiting a pair of light rings, one of which is stable. We have evolved these objects for $\gtrsim$1000 light crossing times, finding no evidence of an instability due to the appearance of the stable light ring, as previously conjectured Keir (2016); Cardoso et al. (2014) and reported for spinning BSs Cunha et al. (2023). We observe excellent agreement between the axisymmetric/3D evolutions and those in spherical symmetry, suggesting that the inclusion of nonspherical degrees of freedom does not significantly affect the stability of these models. Taking into account also the agreement of our numerical results with perturbative predictions for stability and fundamental radial oscillation frequencies, we interpret our findings as strong numerical evidence in favor of the stability of this class of BH mimickers.
@@
As a cautionary note, we add that poor choices of numerical ingredients, in particular the constraint-damping parameter $\kappa_{1}$ of the CCZ4 formulation, can result in seemingly plausible instabilities of our thin-shell BSs. With $M\kappa_{1}\gg 0.1$, for example, both models S06A044 and S08A06 collapse to a BH; the erroneous nature of these simulations reveals itself only through large and growing constraint violations and complete lack of convergence.
@@
If the presence of the stable light ring is not sufficient to destroy these solitonic boson stars, this family of spherically symmetric models, i.e. with quantized spin zero, as well as similar UCOs, can be considered as viable black-hole mimickers again. The thin-shell solutions under consideration here can mimic the phenomenology of black holes in several different ways, as for example in the GW emission from binaries Evstafyeva et al. (2024) or the creation of BH like shadows in VLBI images Rosa et al. (2023). Figure 6 illustrates how the S06A044 model could resemble a Schwarzschild BH with the same ADM mass in the presence of an equatorial, optically thin accretion disk. The image of the BS clearly displays additional structure toward the center of the image, which is a result of the absence of a horizon as well as the pair of light rings. The outermost direct emission band, as well as the first lensed photon ring are nearly indistinguishable for both images.
@@
Figure 6: Side-by-side comparison of ray-traced images of (left) a Schwarzschild BH and (right) the S06_A044 model of the same ADM mass. The color scale indicates the observed brightness intensity, in arbitrary units, from a simple geometrically thin disk located in the equatorial plane. Both panels also provide a close-up of the photon ring region, exhibiting their fractal structure. Images are obtained with FOORT Mayerson ; Mayerson et al. .
@@
Acknowledgments—G.A.M. is supported by the Cambridge Trust at the University of Cambridge. S.S. is supported by the Centre for Doctoral Training at the University of Cambridge funded through STFC. T.E. acknowledges the Perimeter Institute for Theoretical Physics, supported by the Government of Canada through the Department of Innovation, Science and Economic Development and by the Province of Ontario through the Ministry of Colleges and Universities. This work has been supported by STFC Research Grant No. ST/V005669/1. We acknowledge support by the NSF Grants No. PHY-090003, No. PHY-1626190, and No. PHY-2110594; DiRAC projects No. ACTP284 and No. ACTP238; STFC capital Grants No. ST/P002307/1, No. ST/R002452/1, No. ST/I006285/1, and No. ST/V005618/1; and STFC operations Grant No. ST/R00689X/1. Computations were done on the CSD3 and Fawcett (Cambridge), Cosma (Durham), Niagara (Toronto), Narval (Montreal), Stampede2 (TACC) and Expanse (SDSC) clusters.
Supplemental Material
I Convergence Analysis
@@
Light rings: We check convergence of the computed light ring (LR) radii for the S08A06 model by comparing results obtained at resolutions $\mu h\in\{1/12,1/16,1/20\}$ and display in Fig. S1 the resulting areal radii of the stable LR as functions of time. At low resolution the LR radius exhibits a linear drift, much like what was observed for the central amplitude and Kretschmann scalar (cf. Fig. 2 in the main text). Again, this drift converges away at 5th order, and the Richardson extrapolated values, obtained from the two highest-resolution runs and displayed in Fig. S2, oscillate around constant values.
@@
Figure S1: Evolution of the areal radius of the stable LR for the S08A06 model obtained with the E3D code using different resolutions indicated in the legend. The differences between the curves yield convergence at 5th order.
@@
Figure S2: Richardson extrapolation (5th order) of the LR areal radii of the S08A06 model based on the runs at resolutions $\mu h=1/16,1/20$. The bottom panel shows the Richardson extrapolated evolution of the potential difference between the extrema of the AEP.
@@
In Fig. S2 we see that the oscillation amplitude of the stable LR is slightly larger than that of the unstable LR. This effect is more pronounced for the S06A044 model displayed in Fig. 4 of the main text. The reason in both cases is that the unstable light ring is - to good approximation - outside the thin shell, fixing its location close to the Schwarzschild value by Birkhoff’s theorem. The stable light ring, on the other hand, is located in the interior of the star, and therefore affected more strongly by the perturbation.
@@
Figure S3: The left (right) panels illustrate convergence results obtained from Axi simulations of the S08A06 model for the scalar field amplitude (volume-weighted $L_{2}$ norms of the Hamiltonian and momentum constraints). We amplify the differences and constraint magnitudes by factors $Q_{n}$ as expected for $n^{\rm th}$-order convergence.
2D code: We next discuss the convergence results for the S08A06 model obtained in axisymmetry with the Axi code. Using $\mu h=1/8$, $\mu h=1/12$ and $\mu h=1/16$ on the innermost refinement level for the low, medium and high resolutions, respectively, we obtain convergence between fourth and fifth order for the scalar field amplitude; cf. top panel of Fig. S3. In the bottom panels of the figure, we also monitor the volume-weighted $L_{2}$ norms of the constraint violations. We find first-order convergence in the Hamiltonian constraint at very early times which gradually transitions to fourth order at $\mu t\approx 20000$ where it persists throughout the remainder of the simulation. The momentum constraint . . $\mathcal{M}\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=(\mathcal{M}_{x}^{2}+\mathcal{M}_{y}^{2})^{1/2}$ on the other hand rapidly reaches convergence between fourth and fifth order.
II Additional results for $\sigma_{0}=0.06$
@@
In this section we present results for the $\sigma_{0}=0.06$ family of models, supplementing those presented for $\sigma_{0}=0.08$ in the main text. In Fig. S4, we compare the power spectrum of central scalar-amplitude oscillations to the radial oscillation frequency, computed with perturbation theory. In this case the larger error in our initial data causes the system to migrate early on to a boson-star (BS) model with slightly smaller central scalar amplitude than the initial value $A_{0}=0.044$. Following Ref. Kain (2021) we take this effect into account by computing the radial frequency for a BS model with an amplitude averaged over several oscillation cycles, $\langle A_{0}\rangle=0.04394$. Once again, we find very good agreement between results from the perturbative analysis and the codes enforcing different symmetry assumptions.
@@
In Fig. S5, we show timescales of instability for a sample of BS models with $\sigma_{0}=0.06$. As before, we use the time at which the absolute change in the central amplitude $|\Delta A_{0}(t)|$ exceeds $10\%$ of the initial central amplitude $\Delta A_{0}(0).$ Once again, we find that the predictions of the linear radial perturbation theory are borne out by all codes regardless of symmetry assumptions. Models on the first unstable branch either migrate or disperse depending on the sign of the binding energy, and models on the second unstable branch always collapse to a BH. We emphasize that the specific time values depend on factors such as resolution, gauge, evolution method and numerical parameter choices or mesh refinement; values determined with different codes therefore exhibit some variation but always agree on stability or instability. Note also that for some values of $A_{0}$, multiple BS models exist. In those cases, the two lower-mass models are always stable while the heaviest BS may be stable or not, in accordance with the predictions of linear theory.
Figure S4: Power spectrum of $\Delta A_{0}(t)$ for the BS model S06A044; the red line marks the fundamental radial mode frequency $\chi_{0}$. Units on the vertical axis are arbitrary.
III Radial Oscillation Frequencies
@@
In this section we discuss in detail the method used to compute fundamental radial oscillation frequencies for solitonic BSs. Here, we use overdots to denote time derivatives and primes for radial derivatives.
We begin with a spherically symmetric metric ansatz,
$\mathrm{d}s^{2}=-\alpha^{2}(t,r)\mathrm{d}t^{2}+X^{2}(t,r)\mathrm{d}r^{2}+r^{2}\mathrm{d}\Omega_{2}^{2}.$ (S1)
Decomposing the complex scalar field into two real fields, $\varphi(t,r)=\varphi_{1}+i\varphi_{2}$, the Einstein-Klein-Gordon equations become,
$\frac{X^{\prime}}{X}=\,$ $\frac{1-X^{2}}{2r}+2\pi rX^{2}\left(\frac{|\dot{\varphi}|^{2}}{\alpha^{2}}+\frac{|\varphi^{\prime}|^{2}}{X^{2}}+V\right),$ (S2)
$\frac{\alpha^{\prime}}{\alpha}=\,$ $-\frac{1-X^{2}}{2r}+2\pi rX^{2}\left(\frac{|\dot{\varphi}|^{2}}{\alpha^{2}}+\frac{|\varphi^{\prime}|^{2}}{X^{2}}-V\right),$ (S3)
$\varphi_{i}^{\prime\prime}=\,$ $\left(\frac{X^{\prime}}{X}-\frac{\alpha^{\prime}}{\alpha}-\frac{2}{r}\right)\varphi_{i}^{\prime}+\frac{X^{2}\dot{\alpha}+X\alpha\dot{X}}{\alpha^{3}}\dot{\varphi}_{i}+\frac{X^{2}}{\alpha^{2}}\ddot{\varphi}_{i},$ (S4)
$\alpha^{\prime\prime}=\,$ $\frac{X\ddot{X}}{\alpha}-\frac{X\dot{X}\dot{\alpha}}{\alpha^{2}}+\frac{X^{\prime}\alpha^{\prime}}{X}-\frac{\alpha^{\prime}}{r}+\frac{\alpha X^{\prime}}{rX}$ (S5)
@@
We solve Eqs. (S2)-(S4)
Figure S5: Instability timescales as a function of $A_{0}$ for solitonic BSs with $\sigma_{0}=0.06.$ The vertical lines mark the transition between stable and unstable branches as predicted by linear theory. The marker style indicates the dynamical fate while colour indicates the code used: black for Sph, lime green for Axi, orange for E3D. BS models whose time evolution shows no sign of instability are represented with arrows at the top of the figure.
using the ansatz $\varphi(t,r)=A(r)e^{i\omega t}$ to obtain a background solution henceforth denoted by $\{A_{b}(r),\alpha_{b}(r),X_{b}(r),\omega\}$. The background scalar-field amplitude in particular satisfies
$A_{b}^{\prime\prime}=-\frac{\omega^{2}X_{b}^{2}}{\alpha_{b}^{2}}A_{b}+A_{b}^{\prime}\left(\frac{X^{\prime}_{b}}{X_{b}}-\frac{\alpha^{\prime}_{b}}{\alpha_{b}}-\frac{2}{r}\right)+X_{b}^{2}A_{b}V^{\prime}(A_{b}^{2})\,$,\,. (S6)
We perturb the BS background by introducing the perturbations $\delta\psi_{1}$, $\delta\psi_{2}$, $\delta\nu$ and $\delta\lambda$ according to
$\varphi$ $=e^{i\omega t}\left(\psi_{1}+i\psi_{2}\right)\,$,
$\psi_{1}$ $=A_{b}(1+\delta\psi_{1}),\qquad$ $\psi_{2}$ $=A_{b}\delta\psi_{2}\,$,
$\alpha$ $=\alpha_{b}\left(1+\frac{1}{2}\delta\nu\right),\qquad\$$X$ $=X_{b}\left(1+\frac{1}{2}\delta\lambda\right).$ (S7)
Substituting this expansion into Eqs. (S2), (S4), (S5), taking the difference between Eqs. (S3) and (S2), and linearizing, we obtain
@@
$\delta\psi_{1}^{\prime\prime}=\,$ $X_{b}^{2}\left(1-\frac{\omega^{2}}{\alpha_{b}^{2}}-\frac{8A_{b}^{2}}{\sigma_{0}^{2}}+\frac{12A_{b}^{4}}{\sigma_{0}^{4}}\right)\delta\lambda+\frac{X_{b}^{2}}{\alpha_{b}^{2}}\delta\ddot{\psi}_{1}+\frac{16A_{b}^{2}X_{b}^{2}}{\sigma_{0}^{2}}\left(\frac{3A_{b}^{2}}{\sigma_{0}^{2}}-1\right)\delta\psi_{1}+\frac{A_{b}^{\prime}}{2A_{b}}\left(\delta\lambda^{\prime}+4\delta\psi_{1}^{\prime}-\delta\nu^{\prime}\right)$
$+\frac{\omega X_{b}^{2}}{\alpha_{b}^{2}}\left(2\delta\dot{\psi}_{2}-\omega\delta\nu\right)+\left(\frac{X_{b}^{\prime}}{X_{b}}-\frac{\alpha_{b}^{\prime}}{\alpha_{b}}-\frac{2}{r}\right)\delta\psi_{1}^{\prime}\,$, (S8)
@@
$\delta\lambda^{\prime}=\,$ $\left(\frac{2X_{b}^{\prime}}{X_{b}}-\frac{1}{r}-4\pi r(A_{b}^{\prime})^{2}\right)\delta\lambda+8\pi rX_{b}^{2}\left(A_{b}^{2}+\frac{(A_{b}^{\prime})^{2}}{X_{b}^{2}}+\frac{\omega^{2}A_{b}^{2}}{\alpha_{b}^{2}}+\frac{12A_{b}^{6}}{\sigma_{0}^{4}}-\frac{6A_{b}^{4}}{\sigma_{0}^{2}}\right)\delta\psi_{1}$
$+\frac{4\pi r\omega A_{b}^{2}X_{b}^{2}}{\alpha_{b}^{2}}\left(\omega\delta\nu+2\delta\dot{\psi}_{2}\right)+8\pi rA_{b}^{\prime}A_{b}\delta\psi_{1}^{\prime}.$ (S9)
@@
$\delta\ddot{\lambda}=\,$ $\frac{2\alpha_{b}}{X_{b}^{2}}\left[\frac{\alpha_{b}X_{b}^{\prime}}{rX_{b}}+\alpha_{b}^{\prime}\left(\frac{X_{b}^{\prime}}{X_{b}}-\frac{1}{r}\right)-\alpha_{b}^{\prime\prime}-4\pi\alpha_{b}(A_{b}^{\prime})^{2}\right]\delta\lambda-\frac{16\pi\alpha_{b}^{2}}{\omega X_{b}}\left[\frac{(A_{b}^{\prime})^{2}}{X_{b}^{2}}+A_{b}^{2}-\frac{8A_{b}^{4}}{\sigma_{0}^{2}}+\frac{12A_{b}^{6}}{\sigma_{0}^{4}}-\omega^{2}A_{b}^{2}\right]\delta\psi_{1}$
@@
$+8\pi\omega A_{b}^{2}\left(\omega\delta\nu-2\delta\dot{\psi}_{2}\right)+\frac{1}{rX_{b}}\left(\frac{r\alpha_{b}^{2}\delta\nu^{\prime}}{X_{b}}\right)^{\prime}+\frac{16\pi\alpha_{b}^{2}A_{b}^{\prime}A_{b}}{X_{b}^{2}}\delta\psi_{1}^{\prime}-\frac{1}{rX_{b}^{2}}\left(\alpha_{b}^{2}+r\alpha_{b}\alpha_{b}^{\prime}\right)\delta\lambda^{\prime}.$ (S10)
@@
$\delta\lambda^{\prime}-\delta\nu^{\prime}=\,$ $2\left(\frac{X_{b}^{\prime}}{X_{b}}-\frac{\alpha_{b}^{\prime}}{\alpha_{b}}-\frac{1}{r}\right)+16\pi rA_{b}^{2}X_{b}^{2}\left(1-\frac{8A_{b}^{2}}{\sigma_{0}^{2}}+\frac{12A_{b}^{4}}{\sigma_{0}^{4}}\right).$ (S11)
@@
Next, we make the harmonic ansatz $\delta\psi_{1}(t,r)=e^{\mathrm{i}\mkern 1.0mu\chi t}f(r),$ $\delta\lambda(t,r)=e^{\mathrm{i}\mkern 1.0mu\chi t}g(r)$ with $\chi\in\mathbb{C}$ to be determined. The resulting system of two pulsation equations is self-adjoint, so that, in analogy to the BS frequencies $\omega$, the radial oscillation frequencies $\chi$ form an infinite discrete sequence $\chi_{0}^{2}<\chi_{1}^{2}<$..., where the solution with eigenvalue $\chi_{n}^{2}$ is regular with $n$ zero crossings in $f$ and $g$. A negative $\chi_{0}^{2}$ corresponds to a radial instability, while $\chi_{0}^{2}>0$ guarantees (linear) radial stability.